Assume that a sample is used to estimate a population proportion p. Find the 99% confidence interval for a sample of size 305 with 168 successes. Enter your answer as an open-interval (i.e., parentheses) using decimals (not percents) accurate to three decimal places.

C.I. =
Answer #1

Solution :

Given that,

Point estimate = sample proportion = \hat p = x / n = 168 / 305 = 0.551

1 - \hat p = 1 - 0.551 = 0.449

Z\alpha/2 = 2.576

Margin of error = E = Z\alpha / 2 * \sqrt((\hat p * (1 - \hat p)) / n)

= 1.645 * (\sqrt((0.551 * 0.449) / 305)

= 0.073

A 99% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.551 - 0.073 < p < 0.551 + 0.073

0.478 < p < 0.624

C.I. = 0.478 , 0.624